OpenStudy (anonymous):

what fraction is not equal to 4/12?

OpenStudy (karatechopper):

Got some choices?

OpenStudy (anonymous):

8/24 1/3 1/4 32/96

OpenStudy (anonymous):

1/4 is not equal to it 32/96 simplified = 1/3 8/24 simplified = 1/3 and of course 1/3=1/3

OpenStudy (karatechopper):

^

OpenStudy (anonymous):

i just tryed that and i got it wrong

OpenStudy (anonymous):

0.o

OpenStudy (anonymous):

sorry I read the question wrong..

OpenStudy (anonymous):

skyler 1/4 is not equal to 1/3

OpenStudy (anonymous):

it is cool